# The prescribed metric on the boundary of convex subsets of anti-de Sitter space with a quasi-circle as ideal boundary

### v1

@@ 5. Approximation
To do that, we will use curvatures. We will approximate $K_{h}$, the curvature of $h$, by smooth functions $K_{n}$ invariant under $\rho_{n}$. Then, we will show that each $K_{n}$ is the curvature of a metric $h_{n}$ which is invariant under $\rho_{n}$. Finally, we will show that the metric $h_{n}$ converges to $h$ smoothly on compact subsets of $\mathbb{D}$. Note that for us it is important to have bounded derivatives with bounds that do not depend on $n$.
@@
Let $K:\mathbb{D}\to\left(-\frac{1}{\epsilon},-1-\epsilon\right)$ be a smooth function such that any derivative of it at order $p$ is bounded by some $M_{p}>0$ uniformly on $\mathbb{D}$. Let $\rho_{n}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$ be a sequence of Fuchsian representations that have injectivity radius growing to $\infty$. Then there exists a sequence of smooth functions $K_{n}:\mathbb{D}\to\left(-\frac{1}{\epsilon^{\prime}},-1-\epsilon^{\prime}\right)$, such that each $K_{n}$ is $\rho_{n}$-equivariant, $K_{n}$ converges $C^{\infty}$ on compact subsets to $K$, and each derivative of order $p$ of $K_{n}$ is bounded on the disc $\mathbb{D}$ with respect to the hyperbolic metric by some $M^{\prime}_{p}$, where $M^{\prime}_{p}$ does not depend on $n$.
@@
Let $\phi:\mathbb{R}\to\left[0,1\right]$ be a smooth function such that $\phi\mid_{\left[0,1\right]}=1$ and $\phi(x)=0,\forall x\in\mathbb{R}\setminus\left[0,2\right]$ (see Urysohn smooth lemma). We define the function $k^{\prime}_{n}$ to be
$K^{\prime}_{n}=\left\{\begin{matrix}K\ \textit{on}\ B(o,r_{n})\\ A_{n}\ \textit{on}\ B(o,2r_{n})\setminus B(o,r_{n})\\ -\frac{1}{\epsilon}\ \textit{on}\ D_{n}\setminus B(o,2r_{n})\end{matrix}\right.$
where $A_{n}=\phi(\frac{d_{\mathbb{H}^{2}}(o,x)}{r_{n}})K+(1-\phi(\frac{d_{\mathbb{H}^{2}}(o,x)}{r_{n}}))(-\frac{1}{\epsilon})$.
Note that that $K^{\prime}_{n}$ has values in the interval $\left[-\frac{1}{\epsilon},-1-\epsilon\right]$. Also note that the derivatives of $K^{\prime}_{n}$ depend only on $K$ and $\phi$, $d_{\mathbb{H}^{2}}(x,o)$ and $\frac{1}{r_{n}}$ which can be assumed arbitrary small. Then we deduce that any derivative of $K^{\prime}_{n}$ of order $\alpha$ is bounded (with respect to the hyperbolic metric) by a constant $M^{\prime}_{\alpha}$ that does not depend on $n$ (it depends on the function $\phi$ and on the bound of the derivative of $h$ of order $\alpha$) .
@@
We will use the next two theorems
Theorem 5.3.
[Duc91] Let $k:\mathbb{D}\to\mathbb{R}_{-}$ be a $C^{\infty}$ function, then there exists a unique complete metric $h$ on $\mathbb{D}$ which is conformal to $\left|dz\right|^{2}$ and has curvature equal to $k$.
and
@@
If $k_{n}$ converge uniformly $C^{\infty}$ on compact subsets to $k$, then $h_{n}$ converge uniformly $C^{\infty}$ on compact subsets to $h$.
@@
In this proof we will use euclidean norms, and we will use the fact that the hyperbolic metric and the euclidean metric are bi-Lipchitz on compact subsets. We denote
[…]
and we denote
$\left\|f\right\|_{C^{k,\alpha}(B)}:=\sum_{\left|\beta\right|\leq k}\left\|\partial^{\beta}f\right\|_{\infty}+\sum_{\left|\beta\right|=k}\left[\partial^{\beta}f\right]_{\alpha}$
@@
where $\left[\partial^{\beta}f\right]_{\alpha}:=\sup\limits_{0<R<1,z\in B}R^{-\alpha}\sup\left\{\left|u(x)-u(y)\right|;x,y\in B(z,R)\cap B\right\}$ (here, by $B(z,R)$, we mean the euclidean ball).
@@
Recall that at a point $p$ of the disc $\mathbb{D}$, the hyperbolic metric $h_{-1}$ satisfies $h_{-1}:=\frac{4\left|dz\right|^{2}}{(1-\left\|p\right\|^{2})^{2}}$. In particular, the hyperbolic metric and the euclidean metric are bi-Lipschitz to each other on compact subsets of $\mathbb{D}$ with a bi-Lipschitz constant that depends on that compact subset.
@@
We will show that if $K$, the curvature of a complete conformal metric $h$ on $\mathbb{D}$, belongs to an interval of the form $(-\frac{1}{\epsilon},-1-\epsilon)$ and every derivative of $K$ or order $p$ is bounded by $M_{p}>0$, than all the partial derivatives of the metric at any order are bounded by constants that depend only on $M_{p}$ and $\epsilon$.
@@
Fix $r>0$. Denote by $B(x,r)$ the hyperbolic ball of center $x$ and radius $r$. The metric $h=e^{2u}h_{-1}$, where $u$ is a smooth function and $h_{-1}$ is the hyperbolic metric. Note that (see [Tro91]) the function $u$ is bounded on the disc because the curvature of $h$ is negative and bounded.
@@
Up to apply a hyperbolic isometry that sends the ball $B(x,r)$ to the hyperbolic ball $B(o,r)$, we can always remain in the hyperbolic disc $B=B(o,r)$ which is centered at $o$ and have hyperbolic radius $r$. Let $g$ be a hyperbolic isometry, then we get $g^{*}(h)=e^{u^{*}}h_{-1}$, where $u^{*}=u\circ g$, in particular $u^{*}$ still bounded on $\mathbb{D}$.
@@
Since the hyperbolic metric and the euclidean metric are bi-Lipchitz with a bi-Lipchitz constant that depends only on $r$, we find that all the derivatives of $K$ restcited to $B$ are bounded with respect to the euclidean metric at any order. We will show that this implies that the derivatives of $u^{*}$ are bounded by constants that depend only on the bounds of the derivatives of $K$. Again, since the euclidean metric and the hyperbolic metric are bi-Lipchitz with a constant that depends only on $r$ (that we can control), then the derivatives of $u^{*}$ will be bounded by the hyperbolic metrics and the bounds depend only on $r$ and the hyperbolic bounds of the derivatives of $K$. When we apply an isometry to go back to the ball $B(x,r)$, we find that all the derivatives of $u$ are bounded on $B(x,r)$ with respect to the hyperbolic metric, and the bounds depend only on $\epsilon$ and the bounds of $K$.
@@
Now we restrict our self to the hyperbolic ball $B(0,r)$. Recall the equation $e^{2u^{*}}K^{*}+1=-\Delta u^{*}$, where $h^{*}=e^{2u^{*}}h_{1}$ and $\Delta u^{*}$ is the hyperbolic Laplacian and $K^{*}$ is the curvature of $h^{*}$. By [Nic20, Theorem 10.3.1], for any $k\in\mathbb{N}$ and $p>1$ there exists a constant $C_{k,p}$ such that the following inequality holds
$\left\|u^{*}\right\|_{k+2,p}\leq C_{k,p}(\left\|\Delta u^{*}\right\|_{k,p}+\left\|u^{*}\right\|_{p})$
@@
Note that since $u$ is uniformly bounded with bounds that depend only $\epsilon$, and since $k^{*}$ and all its derivatives are uniformly bounded at any order, it follows that by applying induction on the equality $e^{2u^{*}}K^{*}+1=-\Delta u^{*}$ we get the existence of constants $M^{\prime}_{k,p}$ such that $\left\|u^{*}\right\|_{k,p}$ is bounded for any $k,p$ by $M^{\prime}_{k,p}$.
@@
By Morrey inequality there exist constant $C^{\prime}_{k,p}$ such that
$\left\|u^{*}\right\|_{C^{k-\gamma,\gamma-\frac{2}{p}}}\leq C^{\prime}_{k,p}\left\|u^{*}\right\|_{k,p}$
@@
In particular the lemma will follows because the metric $h_{n}$ will converge uniformly $C^{\infty}$ on compact subsets up to extracting a subsequence (because their derivatives are uniformly bounded at any order with respect the hyperbolic metric, this follows from the arguments above). Since the curvatures of $h_{n}$ converge to the curvature of $h$, we obtain that $h_{n}$ converge (up to extracting) to $h$.
Proposition 5.6.
Let $S_{n}$ be a sequence of closed surfaces having genus $g_{n}$ converging to $\infty$. Let $\rho_{n}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$ be a sequence of Fuchsian representations that have injectivity radius converging to $\infty$. Let $h$ be a complete, conformal Riemannian metric on the disc $\mathbb{D}$ that has curvature in $(-\frac{1}{\epsilon},-1-\epsilon)$ for some $\epsilon>0$, and any derivative of $h$ of order $p$ is bounded by some $M_{p}$. Then there exists a sequence $h_{n}$ of complete, conformal Riemannian metrics that have curvature in $(-\frac{1}{\epsilon^{\prime}},-1-\epsilon^{\prime})$, for some $\epsilon^{\prime}>0$, each $h_{n}$ is $\rho_{n}$-invariant, and $h_{n}$ converges $C^{\infty}$ on compact subsets of $\mathbb{D}$ to $h$, and any derivative of $h_{n}$ of order $p$ is bounded by some $M^{\prime}_{p}$, where $M^{\prime}_{p}$ does not depend on $n$.
@@
By Lemma 5.2 we construct $K_{n}:\mathbb{D}\to(-\frac{1}{\epsilon^{\prime}},-1-\epsilon^{\prime})$, a sequence of $\rho_{n}$ invariant functions that converge $C^{\infty}$ on compact subsets to $K_{h}$, where $K_{h}$ is the curvature of $h$ (recall that $K_{n}$ have uniformly bounded derivatives). By Theorem 5.4, for any $n$ there is a unique complete conformal metric $h_{n}$ on the disc which has curvature equal to $h_{n}$, and which is $\rho_{n}$ invariant. Since $K_{n}$ the curvatures of the metrics $h_{n}$ converge smoothly to $K_{h}$, Lemma 5.5 implies that the metrics $h_{n}$ converge, up to extracting a subsequence, to $h$. The derivatives of $h_{n}$ at any order are all uniformly bounded (independently on $n$) by Lemma 5.5. ∎
@@ 6. The proof of the main theorem
In this section we will prove the main theorem. Before proceeding with the proof we need to recall the following statements.
@@
Now we proceed the proof of our main theorem
@@
Let $h^{\pm}$ be two complete conformal Riemannian metrics on the disk $\mathbb{D}$ that have curvature in $(-\frac{1}{\epsilon},-1-\epsilon)$ for some $\epsilon>0$, and each derivative of $h^{\pm}$ of order $p$ is bounded by $M_{p}>0$ (with respect to the hyperbolic metric). Let $f$ be a normalized quasi-symmetric map. There exists a normalized globally hyperbolic convex subset $\Omega$, and normalized isometries $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ such that $f=\Phi_{\Omega}$.
@@
By Proposition 6.2 there are quasi-Fuchsian representations $\rho_{n}^{+},\rho_{n}^{-}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$, and $f_{n}$ a sequence of $\rho_{n}^{+},\rho_{n}^{-}$ equivariant quasi-symmetric maps that converge in the $C^{0}$ topology to $f$. Also by Proposition 5.1 there is a $\rho_{n}^{+}$ (resp $\rho_{n}^{-}$) metrics $h_{n}^{+}$ (resp $h_{n}^{-}$) that converge $C^{\infty}$ on compact subsets of $\mathbb{D}$ to $h^{+}$ (resp $h^{-}$). Also by Proposition 5.1, the metrics $h^{\pm}_{n}$ have bounded uniformly derivatives (that is the bounds depend only on the order of the and do not depend on $n$ or the point on the $\mathbb{D}$).
@@
The globally hyperbolic manifold $M_{n}$ lifts to a globally hyperbolic convex subset $\Omega_{n}$, up to normalizing, we have isometries $V^{\pm}_{n}:(\mathbb{D},h^{\pm}_{n})\to\partial^{\pm}\Omega_{n}$ such that $\partial V_{n}^{\pm}(p)=(p,p)$ for $p=0,1,\infty$. Since there is a unique $\rho_{n}^{+},\rho_{n}^{-}$ equivariant quasi-symmetric map, we have that $\Phi_{\Omega_{n}}=f_{n}$.
By Proposition 6.2, the maps $f_{n}$ are uniformly quasi-symmetric. Then there is $k^{\prime}>1$ such that $\partial_{\infty}\Omega_{n}$ is a $k^{\prime}$ quasi-circle for any $n$. This implies that $\Omega_{n}$ converge in the Hausdorff topology to a globally hyperbolic convex subset $\Omega$ Then by Proposition 4.12, the isometries $V_{n}^{\pm}:(\mathbb{D},h_{n}^{\pm})\to\partial^{\pm}\Omega_{n}$ converge to an isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ and the symmetric map $\Phi_{\Omega}=f$. ∎

### v2

@@ 6. Approximation
To do that, we will use curvatures. We will approximate $K_{h}$, the curvature of $h$, by smooth functions $K_{n}$ invariant under $\rho_{n}$. Then, we will show that each $K_{n}$ is the curvature of a metric $h_{n}$ which is invariant under $\rho_{n}$. Finally, we will show that the metrics $h_{n}$ converge to $h$ smoothly on compact subsets of $\mathbb{D}$. Note that for us it is important to have bounded derivatives with bounds that do not depend on $n$.
@@
Let $K:\mathbb{D}\to\left(-\frac{1}{\epsilon},-1-\epsilon\right)$ be a smooth function such that any derivative of it at order $p$ is bounded by some $M_{p}>0$ uniformly on $\mathbb{D}$. Let $\rho_{n}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$ be a sequence of Fuchsian representations that have injectivity radius growing to $\infty$. Then there exists a sequence of smooth functions $K_{n}:\mathbb{D}\to\left(-\frac{1}{\epsilon^{\prime}},-1-\epsilon^{\prime}\right)$, such that each $K_{n}$ is $\rho_{n}$-equivariant, $K_{n}$ converge $C^{\infty}$ on compact subsets to $K$, and each derivative of order $p$ of $K_{n}$ is bounded on the disc $\mathbb{D}$ by some $M^{\prime}_{p}$, where $M^{\prime}_{p}$ does not depend on $n$.
@@
Let $\phi:\mathbb{R}\to\left[0,1\right]$ be a smooth function such that $\phi\mid_{\left(-\infty,1\right]}=1$ and $\phi\mid_{\left[2,+\infty\right)}=0$ (see Urysohn smooth lemma). We define the function $K^{\prime}_{n}$ on $D_{n}$ by :
$K^{\prime}_{n}(x):=\phi\left(\frac{d_{\mathbb{H}^{2}}(o,x)}{r_{n}}\right)K+\left(1-\phi\left(\frac{d_{\mathbb{H}^{2}}(o,x)}{r_{n}}\right)\right)\left(-\frac{1}{\epsilon}\right).$
In particular, note that $K^{\prime}_{n}\mid_{B(o,r_{n})}=K$ and $K^{\prime}_{n}\mid_{D_{n}\setminus B(o,2r_{n})}=-\frac{1}{\epsilon}$.
Note that that $K^{\prime}_{n}$ has values in the interval $\left[-\frac{1}{\epsilon},-1-\epsilon\right]$. Also note that the derivatives of $K^{\prime}_{n}$ depend only on $K$ and its derivatives, the derivatives of $\phi$ on the closed interval $\left[1,2\right]$, and the derivatives of $\frac{d_{\mathbb{H}^{2}}(o,x)}{r_{n}}$ that are uniformly bounded at any order $\alpha$ on a neighborhood of the subset $B(o,2r_{n})\setminus B(o,r_{n})$. Then we deduce that any derivative of $K^{\prime}_{n}$ of order $\alpha$ is bounded by a constant $M^{\prime}_{\alpha}$ that does not depend on $n$ (it depends on the function $\phi$ and on the bound of the derivative of $K$ of order $\alpha$).
@@
We will use the next theorem :
@@
If $K_{n}$ converge uniformly $C^{\infty}$ on compact subsets to $K_{h}$, then $h_{n}$ converge uniformly $C^{\infty}$ on compact subsets to $h$.
@@
In this proof we will use euclidean norms, and we will use the fact that the hyperbolic metric and the euclidean metric are bi-Lipchitz on compact subsets.
We denote :
$\left\|f\right\|_{k,p}:=(\sum_{\alpha\leq k}\int_{B}\left|\partial^{\alpha}f\right|^{p}dx)^{\frac{1}{p}},$
also we denote :
$\left\|f\right\|_{C^{k,\alpha}(B)}:=\sum_{\left|\beta\right|\leq k}\left\|\partial^{\beta}f\right\|_{\infty}+\sum_{\left|\beta\right|=k}\left[\partial^{\beta}f\right]_{\alpha},$
@@
where $\left[\partial^{\beta}f\right]_{\alpha}:=\sup\limits_{0<R<1,z\in B}R^{-\alpha}\sup\left\{\left|f(x)-f(y)\right|;x,y\in D(z,R)\cap B\right\}$ (here, by $D(z,R)$, we mean the euclidean ball).
@@
Recall that at a point $p$ of the disc $\mathbb{D}$, the hyperbolic metric $h_{-1}$ is defined by $h_{-1}:=\frac{4\left|dz\right|^{2}}{(1-\left\|p\right\|^{2})^{2}}$.
@@
We will show that if $K$, the curvature of a complete conformal metric $h$ on $\mathbb{D}$, belongs to an interval of the form $(-\frac{1}{\epsilon},-1-\epsilon)$ and every derivative of $K$ or order $p$ is bounded by $M_{p}>0$, then all the partial derivatives of the metric at any order are bounded by constants that depend only on $M_{p}$ and $\epsilon$.
@@
Fix $r>0$. Denote by $B(x,r)$ the hyperbolic ball of center $x$ and radius $r$. The metric $h$ satisfies $h=e^{2u}h_{-1}$, where $u$ is a smooth function and $h_{-1}$ is the hyperbolic metric. Note that (see [Tro91]) the function $u$ is bounded on the disc because the curvature of $h$ is negative and bounded.
@@
Up to applying a hyperbolic isometry that sends the ball $B(x,r)$ to the hyperbolic ball $B(o,r)$, we may assume that we remain in the hyperbolic ball $B:=B(o,r)$ centered at $o$ with hyperbolic radius $r$. Let $g$ be a hyperbolic isometry, then we get $g^{*}(h)=e^{u^{*}}h_{-1}$, where $u^{*}=u\circ g$, in particular $u^{*}$ still bounded on $\mathbb{D}$.
@@
Since the hyperbolic metric and the euclidean metric are bi-Lipchitz with a bi-Lipchitz constant that depends only on $r$, we find that all the derivatives of $K$ restcited to $B$ are bounded with respect to the euclidean metric at any order. We will show that this implies that the derivatives of $u^{*}$ are bounded by constants that depend only on the bounds of the derivatives of $K$. Again, since the euclidean metric and the hyperbolic metric are bi-Lipchitz with a constant that depends only on $r$ (that we can control), then the derivatives of $u^{*}$ will be bounded by the hyperbolic metrics and the bounds depend only on $r$ and the hyperbolic bounds of the derivatives of $K$. After applying an isometry that maps the ball $B$ back to the ball $B(x,r)$, we find that all derivatives of $u$ are bounded on $B(x,r)$ with respect to the hyperbolic metric. These bounds depend only on $\epsilon$ and the bounds of $K$.
@@
Now we restrict our self to the hyperbolic ball $B(o,r)$. Recall the equation $e^{2u^{*}}K^{*}+1=-\Delta u^{*}$, where $h^{*}=e^{2u^{*}}h_{1}$ and $\Delta u^{*}$ is the hyperbolic Laplacian and $K^{*}$ is the curvature of $h^{*}$. By [Nic20, Theorem 10.3.1], for any $k\in\mathbb{N}$ and $p>1$ there exists a constant $C_{k,p}$ such that the following inequality holds :
$\left\|u^{*}\right\|_{k+2,p}\leq C_{k,p}(\left\|\Delta u^{*}\right\|_{k,p}+\left\|u^{*}\right\|_{p}).$
@@
Note that since $u^{*}$ is uniformly bounded with bounds that depend only $\epsilon$, and since $K^{*}$ and all its derivatives are uniformly bounded at any order, it follows that by applying induction on the equality $e^{2u^{*}}K^{*}+1=-\Delta u^{*}$ we get the existence of constants $M^{\prime}_{k,p}$ such that $\left\|u^{*}\right\|_{k,p}$ is bounded for any $k,p$ by $M^{\prime}_{k,p}$.
@@
By Morrey inequality there exist constant $C^{\prime}_{k,p}$ such that :
$\left\|u^{*}\right\|_{C^{k-\gamma,\gamma-\frac{2}{p}}}\leq C^{\prime}_{k,p}\left\|u^{*}\right\|_{k,p},$
@@
In particular, the lemma will follows because the metrics $h_{n}$ will converge uniformly $C^{\infty}$ on compact subsets up to extracting a subsequence (because their derivatives are uniformly bounded at any order with respect the hyperbolic metric, this follows from the arguments above). Since the curvatures of $h_{n}$ converge to the curvature of $h$, we obtain that $h_{n}$ converge (up to extracting) to $h$.
Now we give a proof for Proposition 6.1.
@@
By Lemma 6.2 we construct $K_{n}:\mathbb{D}\to(-\frac{1}{\epsilon^{\prime}},-1-\epsilon^{\prime})$, a sequence of $\rho_{n}$ invariant functions that converge $C^{\infty}$ on compact subsets to $K_{h}$, where $K_{h}$ is the curvature of $h$ (recall that $K_{n}$ have uniformly bounded derivatives). By Theorem 6.3, for any $n$ there is a unique complete conformal metric $h_{n}$ on the disc which has curvature equal to $K_{n}$, and which is $\rho_{n}$ invariant. Since $K_{n}$ the curvatures of the metrics $h_{n}$ converge smoothly to $K_{h}$, Lemma 6.4 implies that the metrics $h_{n}$ converge, up to extracting a subsequence, to $h$. The derivatives of $h_{n}$ at any order are all uniformly bounded (independently on $n$) by Lemma 6.4. ∎
@@ 7. The proof of the main theorem
Figure 3. If there is a closed hyperbolic surface $S_{n}$, two Fuchsian representations $\rho_{n}^{+}$ and $\rho_{n}^{-}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$, and two complete Riemannian metrics $h^{+}_{n}$ and $h^{-}_{n}$ on $\mathbb{D}$ that have sectional curvature strictly less than 1-1 and are invariant under the action of $\rho_{n}^{+}$ and $\rho_{n}^{-}$ respectively. Then by Theorem 7.1, we can find a globally hyperbolic manifold diffeomorphic to $M_{n}=S_{n}\times[0,1]$ such that the induced metric on $S_{n}\times\left\{0\right\}$ is homotopic to $h^{+}_{n}/\rho_{n}^{+}$, and the induced metric on $S_{n}\times\left\{1\right\}$ is homotopic to $h^{-}_{n}/\rho_{n}^{-}$. The globally hyperbolic manifolds $M_{n}$ lift to globally hyperbolic convex subsets $\Omega_{n}$. There is a unique quasi-symmetric map equivariant under the action of $\rho^{+}_{n}$ and $\rho^{-}_{n}$, the gluing map of $\Omega_{n}$ must be equal to this map. The convex sets $\Omega_{n}$ will converge to the desired $\Omega$ that we want to realize.
In this section we will prove the main theorem (see Figure 3). Before proceeding with the proof we need to recall the following statements.
@@
Now we proceed the proof of our main theorem.
@@
Let $h^{\pm}$ be two complete conformal Riemannian metrics on the disk $\mathbb{D}$ that have curvature in $(-\frac{1}{\epsilon},-1-\epsilon)$ for some $\epsilon>0$, and each derivative of $h^{\pm}$ of order $p$ is bounded by $M_{p}>0$. Let $f$ be a normalized quasi-symmetric map. There exists a normalized globally hyperbolic convex subset $\Omega$, and normalized isometries $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ such that $f=\Phi_{\Omega}$.
@@
By Proposition 7.2, there are quasi-Fuchsian representations $\rho_{n}^{+},\rho_{n}^{-}:\pi_{1}(S_{n})\to\rm{PSL}(2,\mathbb{R})$, and $f_{n}$ a sequence of $\rho_{n}^{+},\rho_{n}^{-}$ equivariant quasi-symmetric maps that converge in the $C^{0}$ topology to $f$. Also by Proposition 6.1 there is a $\rho_{n}^{+}$ (resp $\rho_{n}^{-}$) invariant metrics $h_{n}^{+}$ (resp $h_{n}^{-}$) that converge $C^{\infty}$ on compact subsets of $\mathbb{D}$ to $h^{+}$ (resp $h^{-}$). Also by Proposition 6.1, the metrics $h^{\pm}_{n}$ have uniformly bounded derivatives.
@@
The globally hyperbolic manifold $M_{n}$ lifts to a globally hyperbolic convex subset $\Omega_{n}$. Up to normalizing, we can find isometries $V^{\pm}_{n}:(\mathbb{D},h^{\pm}_{n})\to\partial^{\pm}\Omega_{n}$ such that $\partial V_{n}^{\pm}(p)=(p,p)$ for $p=0,1,\infty$. Since there is a unique $\rho_{n}^{+},\rho_{n}^{-}$ equivariant quasi-symmetric map, it follows that $\Phi_{\Omega_{n}}=f_{n}$.
By Proposition 7.2, the maps $f_{n}$ are uniformly quasi-symmetric. Then there is $k^{\prime}>1$ such that $\partial_{\infty}\Omega_{n}$ is a $k^{\prime}$ quasi-circle for any $n$. This implies that $\Omega_{n}$ converge in the Hausdorff topology to a globally hyperbolic convex subset $\Omega$. Then by Proposition 5.15, the isometries $V_{n}^{\pm}:(\mathbb{D},h_{n}^{\pm})\to\partial^{\pm}\Omega_{n}$ converge to an isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ and the gluing map $\Phi_{\Omega}$ will be equal $f$. ∎